Using the money multiplier 16.66, how much will the money supply change if checkable deposits increase by $300 million?
1
Expert's answer
2019-05-29T09:27:59-0400

m=1/rr

rr=1/16.16=0.062

RR=300*0.062=18.6 m

Money supply change =(300-18.6)*16.16=4547 $million